# CustomListener: Text-guided Responsive Interaction for User-friendly Listening Head Generation

### v1

@@ 4 Experiments
[…]
@@ 4.1 Datasets
Since the existing listener head datasets do not contain the full data required for training our model, we conduct additional data annotations on two datasets:(1) ViCo [43], a popular dataset in listener head generation; (2) RealTalk, proposed by [12], which serves as a database for retrieving listener videos in [12]. Firstly, to learn the relationships between the text prior and listener expressions, we follow the text annotation pipeline in TalkCLIP [24], and conduct text annotation on ViCo and RealTalk. Specifically, we employ [6] to acquire emotion labels, activated AUs as well as their intensities. Additionally, to learn realistic head movements such as nodding to convey agreement and shaking the head to signify disagreement, we use Hopenet [32] to detect head motions. Finally, we generate diverse text prior description for each video segment according to the format: [A person $<$EMOTION$>$ and listens with $<$AU$>$ (and $<$HEAD MOTION$>$)], where $<$EMOTION$>$ denotes the emotional labels, $<$AU$>$ denotes fine-grained expression-related texts based on AU activation intensities and Facial Action Coding System [10], and $<$HEAD MOTION$>$ indicates head nodding/shaking labels. We use “and” to link when there is more than one activated AU.
@@
For ViCo, which contains three parts: train set $\mathcal{D}_{train}$, test set $\mathcal{D}_{test}$, out-of-domain set $\mathcal{D}_{ood}$ (all identities in $\mathcal{D}_{ood}$ have not appeared in $\mathcal{D}_{train}$), we obtain the transcripts of the speaker’s audios by a speech recognition model [29], and get an extended version of ViCo. For RealTalk, to align the dimension of 3DMM coefficients with ViCo, we re-extract facial 3DMM coefficients using [9] and then acquire two subsets: $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$. These two extended datasets are important for future research on text-guided listener head generation task. Notably, we did not use L2L [25] datasets as we cannot obtain the raw videos of L2L [25] and thus cannot perform text annotations on it.
@@ 4.2 Experimental Settings
$\vspace{-0.1cm}L_{d}=E_{x_{0},t,\epsilon_{t}\sim\mathcal{N}(0,1)}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},c,t)\right\|^{2}\right],\vspace{-0.1cm}$ (11)
where $\epsilon_{t}\sim\mathcal{N}(0,1)$ is the noise at step $t$, $c$ denotes conditions, $x_{0}$ is the original listener motions.
@@
The resolution of facial videos is 256x256, and the FPS is 30. For training, we clip the original videos to several 60-frame video segments. We extract 3DMM coefficients from speaker videos and listener videos by [9], which serve as speaker motions and listener motions, respectively. Following MFR-Net [21], we extract 45-dim acoustic features from audios, containing 14-dim MFCC, 28-dim MFCC-Delta, energy, Zero Crossing Rate and loudness. We utilize AdamW optimizer [23] to train our model with a learning rate of $1e^{-4}$.
@@ 4.3 Quantitative Results
It’s non-trivial to measure the realism of generated listener motions. We choose the following metrics based on that a good custom listener should have four characteristics: (1) natural-looking and diverse facial/head motions; (2) highly synchronous with the speaker motions; (3) highly coherent motions in a long video; (4) high image quality after rendering:
@@
$\bullet$ Variation for Diversity (V-D) : Variance across the time series sequence of expression coefficients or pose coefficients.
@@
[…]
@@
Figure 5: Qualitative comparisons with PCH [15], RLHG [43] based on the same speaker and the same listener reference image.
@@
We retrain PCH [15], RLHG [43] and L2L [25] on VICO dataset. It should be noted that, due to the unavailability of source codes from ELP [33] and MFR-Net [21], for MFR-Net [21], we utilize the data from the original paper. For ELP [33], the 3DMM coefficient extraction model used is different from ours, resulting in distinct dimensions of coefficients (e.g., $\beta\in\mathbb{R}^{100T}$ for ELP [33], $\beta\in\mathbb{R}^{64T}$ for ours), thus it is not reasonable to directly compare with evaluation data in ELP [33]. Therefore, we only provide its visual comparisons in the Appendix C. Our quantitative comparisons consist of four parts: 1) precision and diversity of listener motions: as presented in Table 1, for the realism of listener motion generation, our proposed CustomListener achieves lowest FD on both $\mathcal{D}_{test}$ and $\mathcal{D}_{ood}$ subsets in ViCo. Meanwhile, the V-D of our model is comparable to other methods, which indicates that our model can maintain certain diversity when generating controllable listener motions. This may benefit from the probabilistic nature of the diffusion model. 2) speaker-listener synchronization: compared to all existing methods, CustomListener can decrease the average RTLCC and the RWTLCC,
[…]
Method FD $\downarrow$ RTLCC $\downarrow$ RWTLCC $\downarrow$ FID ${}_{\triangle fm}$$\downarrow$ SND $\downarrow$
@@ 4.4 Qualitative Results
In Figure 5, we compare our results with RLHG [43], PCH [15]. It can be seen that our method is capable of generating more realistic listener heads than other baselines. In the left four columns, both RLHG [43] and PCH [15] suffer from unnatural motions (e.g., disorderly opened mouth) and visible facial artifacts. In the right four columns, RLHG [43] still exhibits incorrect motions and PCH [15] can not display negative attitude accurately. While our methods can produce natural listener head that highly aligns with the ground truth conditioned on the given texts. For more detailed image/video format comparisons, please refer to our Appendix C.
@@ 4.5 Ablation Study
In Section 3.2, we design a SDP module to generate dynamic portrait tokens from static one for the purpose of realizing progressive motion shifts that correspond well to the speaker. To validate its effectiveness, we conduct ablation study on Audio-text Responsive Interaction (ATRI) and Speaker Motion-based Weighting (SMW), respectively. The results are shown in Table 3. In addition to the improvement in FD, the improvement in the RTLCC and RWTLCC is more pronounced, which demonstrates that the SDP module is crucial for enhancing speaker-listener synchronization, and each of its components matters.
@@
[…]

### v2

@@ 4 Experiments
Table 1: Comparisons of our model with other methods on $\mathcal{D}_{test}$ and $\mathcal{D}_{ood}$ of ViCo. Bold represents the best. The means we directly refer to data in their paper and denotes we retrain the model. The $\downarrow$ indicates lower is better and the $\uparrow$ indicates higher is better. The values of FD and FID△fm are multiplied by 100. The quantitative results on RealTalk [12] are in Appendix B.
@@ 4.1 Datasets
Since the existing listener head datasets do not contain the full data required for training our model, we conduct additional data annotations on two datasets:(1) ViCo [46], a popular dataset in listener head generation; (2) RealTalk, proposed by [12], which serves as a database for retrieving listener videos in [12]. Firstly, to learn the relationships between the text prior and listener expressions, we follow the text annotation pipeline in TalkCLIP [24], and conduct text annotation on ViCo[46] and RealTalk[12]. Specifically, we employ [6] to acquire emotion labels, activated AUs as well as their intensities. Additionally, to learn realistic head movements such as nodding to convey agreement and shaking the head to signify disagreement, we use Hopenet [34] to detect head motions. Finally, we generate diverse text prior description for each video segment according to the format: [A person $<$EMOTION$>$ and listens with $<$AU$>$ (and $<$HEAD MOTION$>$)], where $<$EMOTION$>$ denotes the emotional labels, $<$AU$>$ denotes fine-grained expression-related texts based on AU activation intensities and Facial Action Coding System [10], and $<$HEAD MOTION$>$ indicates head nodding/shaking labels. We use “and” to link when there is more than one activated AU. Detailed examples can be seen in Appendix A.
@@
For ViCo, which contains three parts: train set $\mathcal{D}_{train}$, test set $\mathcal{D}_{test}$, out-of-domain set $\mathcal{D}_{ood}$ (all identities in $\mathcal{D}_{ood}$ have not appeared in $\mathcal{D}_{train}$), we obtain the transcripts of the speaker’s audios by a speech recognition model [29], and get an extended version of ViCo. For RealTalk, to align the dimension of 3DMM coefficients with ViCo, we re-extract facial 3DMM coefficients using [9] and then obtain two subsets: $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$. These two extended datasets are important for future research on text-guided listener head generation task.
@@ 4.2 Experimental Settings
$\vspace{-0.1cm}L_{d}=E_{x_{0},t,\epsilon_{t}\sim\mathcal{N}(0,I)}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},c,t)\right\|^{2}\right],\vspace{-0.1cm}$ (11)
where $\epsilon_{t}\sim\mathcal{N}(0,\textbf{I})$ is the noise at step $t$, $c$ denotes conditions, $x_{0}$ is the original listener motions.
@@
The resolution of videos is 256$\times$256, and the FPS is 30. For training, we clip the original videos to several 60-frame video segments. We extract 3DMM coefficients from speaker videos and listener videos by [9], which serve as speaker motions and listener motions, respectively. Following MFR-Net [21], we extract 45-dim acoustic features from audio. We utilize AdamW optimizer [23] to train our model with a learning rate of $1e^{-4}$.
@@ 4.3 Quantitative Results
It’s non-trivial to measure the realism of generated listener motions. We choose the following metrics based on that a good custom listener should have four characteristics: (1) natural-looking and diverse facial/head motions; (2) highly synchronous with the speaker motions; (3) highly coherent motions in a long video; (4) high image quality:
@@
$\bullet$ Variation for Diversity (V-D) : Variance across the time series sequence of 3DMM coefficients.
@@
$\bullet$ FID△fm: FID score of differences in 3DMM coefficients between adjacent frames, which indicates the temporal naturalness of facial motions, followed by [44].
@@
We retrain PCH [15], RLHG [46] and L2L [25] on ViCo[46]. It should be noted that, source codes of ELP [35] and MFR-Net [21] is unavailable. For MFR-Net [21], we utilize the data from the original paper. For ELP [35], the 3DMM coefficient extraction model used is different from ours, resulting in distinct dimensions of coefficients (e.g., $\beta\in\mathbb{R}^{100T}$ for ELP [35], $\beta\in\mathbb{R}^{64T}$ for ours), thus it is not reasonable to directly compare with evaluation data in ELP [35]. Therefore, we only provide its visual comparisons in Appendix C. Our quantitative comparisons consist of four parts: 1) precision and diversity of listener motions: as presented in Table 1, for the realism of listener motion generation, our proposed method achieves the lowest FD on both $\mathcal{D}_{test}$ and $\mathcal{D}_{ood}$ in ViCo. Meanwhile, the V-D of our model is comparable to other methods, which indicates that our model can maintain certain diversity when generating controllable listener motions. This may benefit from the probabilistic nature of the diffusion model. 2) speaker-listener synchronization: compared to all existing methods, CustomListener can decrease the RTLCC and the RWTLCC, which justifies the effectiveness of our dynamic portrait tokens. 3) temporal smoothness: in Table 1, the FID△fm and SND of our model outperform other existing methods by a large margin, which validates that our model can mitigate un-smooth transitions. 4) image-level quality: although image quality is not our main focus, we still have a certain performance gain in related metrics, as shown in Table 2. These results indicate that our generated motion is closer to GT motion, thus achieving relatively high image-level quality without special design in rendering.
Figure 5: Qualitative comparisons with PCH [15], RLHG [46] conditioned on the same speaker and the same listener reference image.
4.4 Qualitative Results
Visual Results
Method FD $\downarrow$ RTLCC $\downarrow$ RWTLCC $\downarrow$ FID $\downarrow$ SND $\downarrow$
@@ 4.4 Qualitative Results
In Figure 5, we compare our results with RLHG [46], PCH [15]. It can be seen that our method is capable of generating more realistic listener heads than other baselines. In the left four columns, both RLHG [46] and PCH [15] suffer from unnatural motions (e.g., disorderly opened mouth) and visible facial artifacts. In the right four columns, RLHG [46] still exhibits incorrect motions and PCH [15] can not display negative attitude accurately. While our method can produce natural listener head that highly aligns with the ground truth conditioned on the given texts. More detailed comparisons can be seen in Appendix C.
User Study
We conduct user studies to evaluate from human perceptual perspective. We randomly choose 25 speaker videos from ViCo[46] and RealTalk[12] to generate listeners for each method. 28 people are asked to rate different methods(1-4, 4 is the best) for each video from 4 metrics: similarity to GT, overall naturalness, speaker-listener coordination and video smooth. The average scores of each method are shown in Table 4. Our method outperformed other methods in all aspects.
Method Similarity Overall Speaker-listener Video
to GT Naturalness Coordination Smooth
RLHG [46] 1.28 1.12 1.64 2.04
PCH [15] 2.08 2.40 2.48 2.28
L2L [25] 2.68 2.56 1.92 1.80
Ours 3.96 3.92 3.96 3.88
Table 4: User study results. The best results are highlighted in bold.
@@ 4.5 Ablation Study
In Section 3.2, we design a SDP module to generate dynamic portrait tokens for realizing progressive motion shifts that correspond well to the speaker. To validate its effectiveness, we conduct ablation study on Audio-text Responsive Interaction (ATRI) and Speaker Motion-based Weighting (SMW). The results are shown in Table 3. In addition to the improvement in FD, the improvement in the RTLCC and RWTLCC is more pronounced, which demonstrates the SDP module is crucial for enhancing speaker-listener synchronization, and each of its components matters.
@@
To validate the effectiveness of our Past-guided module (PGM), we train networks without PGM. As presented in Table 3, without PGM, the metrics FID△fm and SND, which are related to the temporal naturalness of motions, increase significantly, indicating a decrease in smoothness. This observation shows the necessity of PGM for mitigating unnatural transitions between video clips. The visual results are shown in Figure 6. Apart from this, our proposed PGM can maintain the consistency of listener’s customized behavioral habits between adjacent clips, as shown in Figure 7. The listener is customized a habit of frowning when thinking, and when segment 2 is guided by “lips mildly parted and lip corner raised slightly”, the frowning habit can be still maintained.
@@ 5 Conclusion
Supplementary Material
Appendix A Implementation Details
A.1 Network Details
We introduced Audio Encoder, Mapping Net, Motion Encoder and Transformer Decoder-based Diffusion Model in Section 3 of our main paper. Due to the page limitation, we show implementation details of these modules in this Appendix.
Audio Encoder
For audio preprocessing, we first extract 45-dim acoustic features from speaker audio, and then encode these acoustic features into audio features. The structure of Audio Encoder is shown in Figure 8, which consists of two convolution layers, a GELU activation function, six 768-dim transformer layers, and a linear layer.
Mapping Net
As discussed in Section 3.2 of the main paper, in order to obtain portrait-related tokens, we further employ a Mapping Net to convert text embeddings from RoBERTa [22] into static portrait tokens, which describes the listener’s static portrait (e.g. expressions, head movements) in the entire video segment. The structure of Mapping Net is shown in Figure 8.
Motion Encoder
We only use a single linear layer to encode the 70-dim noised listener motions to 768-dim motion tokens (denoted with small blue squares in Figure 2 of our main paper).
Transformer Decoder-based Diffusion Model
We employ transformer decoder as our basic denoising module for two reasons: First, compared to U-Net [33] used in traditional diffusion model, transformer [38] can better capture temporal information in listener motion sequences. Furthermore, compared to transformer encoder which only contains self-attention, transformer decoder is capable of facilitating information interaction better between noised listener motion features as well as concatenated dynamic portrait token and motion prior. Specifically, we use an 8-layer transformer decoder, with 8 heads and 2048-dim fully forward layers. Additionally, we concatenate the input motion tokens with diffusion time-step token (denoted with small pink squares in Figure 2 of our main paper) to inject time-step information.
Figure 8: Stucture of Audio Encoder and Mapping Net.
A.2 Inference Details
Long-video Inference
During inference, we generate long-term video in a segment-by-segment way. Specifically, we first sequentially clip the original long video to several 60-frame video segments, containing video frames, audios and speech content. We then input user-customized listener attributes. Subsequently, by incorporating speech content of each segment as well as pre-customized attribute description, text priors of each segment will be produced with the help of GPT. Finally, our proposed CustomListener will generate long-term listener motions clip-by-clip.
Figure 9: Details of Text Prior Generation.
Figure 10: Illustration of Chain-of-thought Prompting.
Text Prior Generation
To generate text prior in specific format like [A person $<$EMOTION$>$ and listens with $<$AU$>$ (and $<$HEAD MOTION$>$)]), inspired by [12], we use chain-of-thought prompting [40] when querying GPT. Specifically, before asking question, we first provide GPT with a few examples showing the desired answer format as well as reasoning process. This strategy can also enable GPT to provide more accurate text prior about facial expressions and head motions while reducing the misunderstanding probability towards the speaker’s content. We present some examples of queries in Figure 9 and Figure 10.
After getting the text prior, we randomly replace the adjectives and adverbs to enhance the diversity of description. For example, for $<$EMOTION$>$, if it is 'happy', we replace it by randomly selecting one from a group $<$looks happy, seems joyful, …$>$. For $<$AU$>$, following Facial Action Coding System, if AU12 (lip corner puller) is activated and level=3, we randomly choose an adj. from $<$raised, pulled, …$>$, an adv. from $<$fully, extremely, …$>$, then combine the above and get [ A person seems joyful and listens with fully raised lip corners.].
[…]
exp angle trans exp pose exp pose exp pose exp pose exp pose
RLHG [46] 20.11 10.32 7.16 0.162 0.365 0.160 0.363 12.42 0.76 4.75 0.09 0.54 0.14
PCH [15] 23.07 13.46 8.04 0.170 0.368 0.163 0.391 12.38 0.81 5.02 0.11 0.29 0.10
L2L [25] 19.01 10.16 7.03 0.158 0.319 0.156 0.315 10.80 0.71 4.67 0.10 0.56 0.40
Ours 17.63 9.30 6.49 0.138 0.218 0.134 0.207 7.70 0.14 4.37 0.07 2.90 1.01
Table 5: Comparisons of our model with other methods on $\mathcal{D}_{test}$ of RealTalk. Bold represents the best. The denotes we retrain the model. The $\downarrow$ indicates lower is better and the $\uparrow$ indicates higher is better. The values of FD and FID△fm are multiplied by 100.
Method SSIM $\uparrow$ CPBD $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
RLHG [46] 0.51 0.29 15.34 30.22
PCH [15] 0.56 0.32 16.13 24.68
L2L [25] 0.56 0.31 15.98 25.68
Ours 0.58 0.34 16.01 23.86
Table 6: Quantitative results with other methods on image quality. The denotes we retrain the model. The best results are highlighted in bold.
Appendix B Quantitative Results on RealTalk
We did not present quantitative results on RealTalk in main paper due to page limitation, thus we give these results in the Appendix B. For a fair comparison, we retrain PCH [15], RLHG [46] and L2L [25] on RealTalk dataset. Due to the unavailability of source codes, we cannot compare with MFR-Net [21] and ELP [35] on RealTalk. Notably, since PCH [15] and RLHG [46] need listener’s attitude label as input, we use [6] to obtain emotion labels of each video clip. Then, we group these emotion labels into three basic attitudes: positive, negative and neutral. As shown in Table 5, we present the FD, RTLCC, RWTLCC, FID△fm, SND and V-D of our model as well as other methods. Our proposed CustomListener achieves best performance on RealTalk across all metrics, which demonstrates the superiority of CustomListener in generating realistic and coherent listener motions that highly synchronous with the speaker’s motions. In Table 6, we also present evaluation results related to image-level quality on RealTalk. Our model achieves optimal results, which indicates CustomListener can generate relatively high-quality listener images without special design in the render module, thus justifies our generated motions are highly precise.
Appendix C Supplementary Visual Results
C.1 Visual Comparisons with Other Methods
In Figure 11, we compared our results with PCH [15], RLHG [46] and ELP [35]. As there is no source code provided by ELP [35], for fairly comparison, we first find the speaker video showed in ELP [35], and then generate listener head based on the speaker video and corresponding description texts instead of an attitude label. Then, we displayed four generated listener frames, which are aligned with the four speaker frames in time. As shown in the visual comparisons, our method has two advantages. Firstly, without any special design in render model, our generated listener images showed less facial artifacts and more closed to the ground truth, which indicates the listener motions produced by CustomListener are highly aligned with the ground-truth motions. Secondly, without a blink modeling like ELP [35], our generated listener can perform blinking action, which demonstrates that the listener motions are natural-looking and photorealistic. More visual comparisons with PCH [15], RLHG [46] and L2L [25] on ViCo and RealTalk are presented in Figure 12 and Figure 13, respectively.
Figure 11: Qualitative comparisons with PCH [15], RLHG [46] and ELP [35].
Figure 12: Qualitative comparisons with other methods on ViCo dataset.
Figure 13: Qualitative comparisons with other methods on RealTalk dataset.
Figure 14: Ablation study of SDP module.
C.2 Visual Results of Ablation
In Section 4.5 of our main paper, we presented visual ablation of PGM module. Due to the page limitation, we show the visual ablation of SDP module in this subsection. As shown in Figure 14, with the help of SDP module, our generated listener motions can not only realize progressive motion changes, but also fluctuate with the speaker’s semantics, intonation and movement amplitude, and thus achieve the speaker-listener coordination.
C.3 Supplementary Video
We provide a supplementary video to present more visual results, which include visual videos on ViCo and RealTalk, more visual comparisons with other methods as well as ablation study of each component. The video can be found in the “Supplementary_Videos.mp4” in the Supplementary Materials. These videos are also displayed on the project page.